OpenStudy (koolkid98):

What is the slope of the line whose equation is 2x – 4y = 10? ? ------ ?

OpenStudy (koolkid98):

@Oleg3321

OpenStudy (oleg3321):

2x - 4y = 10 => Factor out 2 from the left hand side, i.e. 2(x - 2y) = 10 x - 2y = 5 => Set x - 2y = 5 in the slope intercept form: y = mx + c, where m = co-efficient of x = slope, then x - 2y = 5 → y = mx + c - 2y = - x + 5 → y = mx + c 2y = x - 5 → y = mx + c y = x/2 - 5/2 → y = mx + c => From what you have above, m = slope of the line: 2x - 4y = 10 is 1/2

OpenStudy (notapplicable):

Convert to slope intercept form.

OpenStudy (notapplicable):

@Oleg3321 Answers are not allowed..

OpenStudy (koolkid98):

leave him be @NotApplicable

OpenStudy (oleg3321):

are you kidding me? i showed him all the work. you just joined the site so dont tell me what to do lol

OpenStudy (koolkid98):

good job @Oleg3321

OpenStudy (notapplicable):

No, you still gave the answer. You're supposed to guide him..he can still take the answer without looking at the work.

OpenStudy (oleg3321):

well thats not my problem anymore, i showed him the work, if he doesnt want to look at it thats his problem.

OpenStudy (koolkid98):

just STOP fighting

OpenStudy (notapplicable):

No, you have to guide him so he can't take the answer. Make him do some of the work on his own..so he actually understands

OpenStudy (oleg3321):

Whatever im going to leave now, im not gonna get my account suspended because of arguing with you. bye and have a nice day

OpenStudy (notapplicable):

That way, the next time he encounters a problem like this he can do it on his own.

OpenStudy (will.h):

you are supposed to guide him lol he gave him an excellent explanation that's good enough

OpenStudy (notapplicable):

No, it's not.

OpenStudy (koolkid98):

YES IT IS!

OpenStudy (koolkid98):

@NotApplicable

OpenStudy (oleg3321):

3v1 and your still talking lmao

OpenStudy (notapplicable):

@.Sam. @kidrah69

OpenStudy (oleg3321):

yes please bring some mods, lets make something out of nothing

OpenStudy (notapplicable):

I don't care how many people are against me.

OpenStudy (notapplicable):

It doesn't make a difference as long as you're doing wrong..

OpenStudy (koolkid98):

what are you doing im going to posiblly going to get banned

OpenStudy (notapplicable):

You won't get banned smh

OpenStudy (koolkid98):

how do you know? i already had my question closed by mod already @NotApplicable

OpenStudy (oleg3321):

@.Sam.

OpenStudy (koolkid98):

really? i all ready closed it

OpenStudy (notapplicable):

They don't ban for stuff like this unless you keep doing it.

OpenStudy (oleg3321):

nobody did anything except you. first you say i dont know how to do this type of math, then delete the comment after i gave a perfect explenation and get all mad and start making a huge deal out of it. we'll wait till a mod comes and we'll see what they have to say about it smh

OpenStudy (notapplicable):

That was an accidental paste I quoted that from somebody else earlier..my box was glitched and I couldn't click the post button and I was trying to fix it.

OpenStudy (notapplicable):

I accidentally posted that "you don't know how to do this type of math"

OpenStudy (oleg3321):

omg thats the lamest lie ive heard to date lmao

OpenStudy (notapplicable):

It's not a lie..

OpenStudy (koolkid98):

prove it

OpenStudy (oleg3321):

whatever, have a nice day, im not getting my account suspended. peace

OpenStudy (koolkid98):

same here peace out homies

OpenStudy (notapplicable):

http://openstudy.com/study#/updates/57e1f844e4b0dc03c9746463 TheSmartOne said that on this post..but the mods deleted the comments because it was spam..

OpenStudy (koolkid98):

okay? so

OpenStudy (notapplicable):

I was quoting him..and I had that on my clipboard.

OpenStudy (koolkid98):

oh

OpenStudy (koolkid98):

well im going to leave so i dont suspended

OpenStudy (notapplicable):

lol..you kids are scared over nothing

OpenStudy (koolkid98):

i all ready got banned off of peeranwser for this

OpenStudy (koolkid98):

@NotApplicable

OpenStudy (koolkid98):

i dont want to get banned again

OpenStudy (koolkid98):

and it wasnt even my falt

OpenStudy (koolkid98):

both of those mods are offline @NotApplicable

OpenStudy (notapplicable):

PeerAnswer is not OpenStudy smh